OpenStudy (anonymous):

A helicopter flies horizontally at 100 km. it takes 20 seconds for the helicopter to fly from directly overhead to being at an angle of elevation of 60° Find the height of the helicopter above the ground?

OpenStudy (mathstudent55):

You mean it flies at 100 km/h? If so, what distance does it fly in 20 seconds? s = d/t d = st = 100 km/h * 20 s * h/(3600 s) = 0.556 km